### v2

# A Survey on Large Language Model based Human-Agent Systems

Abstract
Recent advances in large language models (LLMs) have sparked growing interest in building fully autonomous agents. However, fully autonomous LLM-based agents still face significant challenges, including limited reliability due to hallucinations, difficulty in handling complex tasks, and substantial safety and ethical risks, all of which limit their feasibility and trustworthiness in real-world applications. To overcome these limitations, LLM-based human-agent systems (LLM-HAS) incorporate human-provided information, feedback, or control into the agent system to enhance system performance, reliability and safety. This paper provides the first comprehensive and structured survey of LLM-HAS. It clarifies fundamental concepts, systematically presents core components shaping these systems, including environment & profiling, human feedback, interaction types, orchestration and communication, explores emerging applications, and discusses unique challenges and opportunities. By consolidating current knowledge and offering a structured overview, we aim to foster further research and innovation in this rapidly evolving interdisciplinary field. Paper lists and resources are available at GitHub repository.
A Survey on Large Language Model based Human-Agent Systems
Henry Peng Zou1, *, Wei-Chieh Huang1, *, Yaozu Wu2, *, Yankai Chen1, †, Chunyu Miao1, Hoang Nguyen1, Yue Zhou1, Weizhi Zhang1, Liancheng Fang1, Langzhou He1, Yangning Li3, Dongyuan Li2, Renhe Jiang2, Xue Liu4,5, Philip S. Yu1, † 1University of Illinois Chicago, 2University of Tokyo, 3Tsinghua University, 4MBZUAI, 5McGill University {pzou3, whuang80, psyu}@uic.edu, yaozuwu279@gmail.com, yankaichen@acm.org
Equal Contribution. Corresponding Author.
1 Introduction
Recent advances in Large Language Models (LLMs) have led to growing enthusiasm for building fully autonomous agent systems that use LLMs as a central engine to perceive environments, make decisions, and execute actions to achieve goals Wang et al. (2024a); Li et al. (2024a). These agents are often equipped with modules for memory, planning, and tool use, aiming to automate complex workflows with minimal human involvement Xie et al. (2024); Xi et al. (2025). However, the pursuit of full autonomy faces critical hurdles. (1) Reliability remains a major concern due to LLMs’ propensity for hallucination, generating plausible but factually incorrect or nonsensical outputs, which undermines trust and can lead to significant errors, especially when actions are chained Gosmar and Dahl (2025); Xu et al. (2024); Glickman and Sharot (2025). (2) Complexity often stalls autonomous agents; they struggle with very complicated tasks requiring deep domain expertise, long multi-step execution, nuanced reasoning, dynamic adaptation, or strict long-context consistency dependencies, as seen in scientific research Feng et al. (2024); Yehudai et al. (2025). (3) Safety and Ethical Risks escalate with autonomy; agents can take unintended harmful actions, amplify societal biases present in training data, or create accountability gaps, particularly in critical decision-making scenarios involving finance, healthcare, or security Mitchell et al. (2025); Deng et al. (2024); Wang et al. (2024c).
Figure 1: Overview of LLM-based Human-Agent Systems (LAM-HAS). LLM-HAS are interactive frameworks where humans actively provide additional information, feedback, or control during interaction with an LLM-powered agent to enhance system performance, reliability, and safety. The system is composed of five core components: Environment & Profiling (including environment settings, and role definitions, goals, and agent capabilities such as planning and memory), Human Feedback (with varying types, timing, and granularity), Interaction Types (collaborative, competitive, cooperative, or mixed), Orchestration (task strategy and temporal synchronization), and Communication (information flow structure and mode).
The persistence of these challenges suggests that full autonomy may be unsuitable for many real-world applications Mitchell et al. (2025); Natarajan et al. (2025) and underscores a crucial insight often overlooked in the drive for pure automation: the indispensable role of human involvement. Humans are frequently needed to provide additional information, essential clarification, or domain knowledge, offer vital feedback and corrections, and exercise necessary oversight and control. These motivate a paradigm shift towards systems explicitly designed for human-agent collaboration: LLM-based Human-Agent Systems (LLM-HAS).
While surveys on LLM-based autonomous agents Wang et al. (2024a); Li et al. (2024a), multi-agent systems Tran et al. (2025); Wu et al. (2025), and specific applications exist Wang et al. (2025b); Peng et al. (2025), a dedicated synthesis focusing specifically on LLM-based human-agent systems is lacking. This survey fills the gap by providing a comprehensive and structured overview of the LLM-HAS. It clarifies the fundamental concepts and systematically presents its core components, emerging applications, and unique challenges and opportunities within this specific niche. To the best of our knowledge, this is still the first survey on LLM-based human-agent systems. We aim to consolidate current knowledge and inspire further research and application in this rapidly evolving interdisciplinary field. We maintain an open-source GitHub repository to provide a sustainable resource complementing our survey paper.
This survey is organized as follows: Section 2 defines and formulates LLM-HAS. Section 3 details the core components shaping the human-agent systems (e.g., human feedback, interaction type, orchestration and communication protocols). Section 4 explores diverse application domains. Section 5 presents open-source implementation frameworks as well as datasets and benchmarks. Section 6 discusses key challenges and future opportunities.
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, anchor=center, align=center, text centered, rectangle, draw=hidden-draw, rounded corners, align=left, minimum width=2em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, font=, [ LLM-based Human-Agent Systems, ver [ Human
Feedback
(Sec 3.2), fill= HumanFeedbackClr, text width= 7em, anchor=center, align=center, font=[ Feedback
Type, fill=HumanFeedbackLight, text width= 5.5em, anchor=center, align=center, font=[ Evaluative, fill=HumanFeedbackLight, text width= 5.5em, anchor=center, align=center, font=[ AgentCoord Pan et al. (2024a), CowPilot Huq et al. (2025),
ConvCodeWorld Han et al. (2025), GDfC Wang et al. (2025c), DigiRL Bai et al. (2024), leaf, fill=HumanFeedbackLight, text width=45em, anchor=center, align=center, text centered, font = ] ] [ Corrective, fill=HumanFeedbackLight, text width= 5.5em, anchor=center, align=center, font=[ Collaborative Gym Shao et al. (2024), ReHAC Feng et al. (2024),
AI Chains Wu et al. (2022), HRC DMP Liu et al. (2024a), BPMN Ait et al. (2024), leaf, fill=HumanFeedbackLight, text width=45em, anchor=center, align=center, text centered, font = ] ] [ Guidance, fill=HumanFeedbackLight, text width= 5.5em, anchor=center, align=center, font=[ FineArena Xu et al. (2025), InteractGen Sun et al. (2024b),
ConvCodeWorld Han et al. (2025), DPT Agent Zhang et al. (2025), Organized Teams Guo et al. (2024b), leaf, fill=HumanFeedbackLight, text width=45em, anchor=center, align=center, text centered, font = ] ] [ Implicit, fill=HumanFeedbackLight, text width= 5.5em, anchor=center, align=center, font=[ Prison Dilemm Jiang et al. (2025), AXIS Lu et al. (2024), AssistantX Sun et al. (2024a),
EmoAgent Qiu et al. (2025), leaf, fill=HumanFeedbackLight, text width=45em, anchor=center, align=center, text centered, font = ] ] ] [ Feedback
Granularity, fill= HumanFeedbackLight, text width= 5.5em, anchor=center, align=center, font=[ Holistic, fill=HumanFeedbackLight, text width= 5.5em, anchor=center, align=center, font=[ Help Feedback Mehta et al. (2024), PARTNR Chang et al. (2024), A2C Tariq et al. (2025),
CoELA Zhang et al. (2024a), DPT Agent Zhang et al. (2025), leaf, fill=HumanFeedbackLight, text width=45em, anchor=center, align=center, text centered, font =] ] [ Segment, fill=HumanFeedbackLight, text width= 5.5em, anchor=center, align=center, font=[ CowPilot Huq et al. (2025), SWEET-RL Zhou et al. (2025), PDFChatAnnotator Tang et al. (2024),
TaPA Wu et al. (2023), leaf, fill=HumanFeedbackLight, text width=45em, anchor=center, align=center, text centered, font = ] ] ] [ Feedback
Phase, fill= HumanFeedbackLight, text width= 5.5em, anchor=center, align=center, font=[ Initial Task, fill=HumanFeedbackLight, text width= 5.5em, anchor=center, align=center, font=[ AgentCoord Pan et al. (2024a), AssistantX Sun et al. (2024a), Hierarchical AgentLiu et al. (2023b), leaf, fill=HumanFeedbackLight, text width=45em, anchor=center, align=center, text centered, font = ] ] [ During Task, fill=HumanFeedbackLight, text width= 5.5em, anchor=center, align=center, font=[ ConvCodeWorld Han et al. (2025), HRC Manipulation Liu et al. (2023a), A2C Tariq et al. (2025), leaf, fill=HumanFeedbackLight, text width=45em, anchor=center, align=center, text centered, font = ] ] [ Post-Task, fill=HumanFeedbackLight, text width= 5.5em, anchor=center, align=center, font=[ GDfC Wang et al. (2025c), SOTOPIA Zhou et al. (2024),DigiRL Bai et al. (2024),
EmoAgent Qiu et al. (2025), leaf, fill=HumanFeedbackLight, text width=45em, anchor=center, align=center, text centered, font = ] ] ] ] [ Interaction
Type
(Sec 3.3), fill=interactClr, text width= 7em, anchor=center, align=center, font=[ Collaboration, fill=interactLight, text width= 6em, anchor=center, align=center, font=[ Supervision, fill=interactLight, text width= 6em, anchor=center, align=center, font=[ EasyLAN Pan et al. (2024b), CowPilot Huq et al. (2025), Hierarchical Agent Liu et al. (2023b),
ConvCodeWorld Han et al. (2025), HRC Manipulation Liu et al. (2023a), leaf, fill=interactLight, text width=44em, anchor=center, align=center, text centered, font = ] ] [ Delegation, fill=interactLight, text width= 6em, anchor=center, align=center, font=[ Collaborative Gym Shao et al. (2024), GDfC Wang et al. (2025c), HRC Manufa Lim et al. (2024),
PaLM-E Driess et al. (2023), DigiRL Bai et al. (2024), leaf, fill=interactLight, text width=44em, anchor=center, align=center, text centered, font = ] ] [ Coordination, fill=interactLight, text width= 6em, anchor=center, align=center, font=[ MTOM Zhang et al. (2024b), Ask Before Plan Zhang et al. (2024c), MindAgent  Gong et al. (2023),
MetaGPT  Hong et al. (2023), leaf, fill=interactLight, text width=44em, anchor=center, align=center, text centered, font = ] ] [ Cooperation, fill=interactLight, text width= 6em, anchor=center, align=center, font=[ CoELA Zhang et al. (2024a), InteractGen Sun et al. (2024b), PARTNR Chang et al. (2024),
MAIH Wang et al. (2024c), leaf, fill=interactLight, text width=44em, anchor=center, align=center, text centered, font = ] ] ] [ Competition, fill=interactLight, text width= 6em, anchor=center, align=center, font=[ SOTOPIA Zhou et al. (2024), leaf, fill=interactLight, text width=51.6em, anchor=center, align=center, text centered, font = ] ] [ Coopetition, fill=interactLight, text width= 6em, anchor=center, align=center, font=[ Prison Dilemm Jiang et al. (2025), SOTOPIA Zhou et al. (2024), leaf, fill=interactLight, text width=51.6em, anchor=center, align=center, text centered, font = ] ] ] [ Orchestration
Paradigm
(Sec 3.4), fill=orchestClr, text width= 7em, anchor=center, align=center, font=[ Task
Strategy, fill=orchestLight, text width= 7.2em, anchor=center, align=center, font=[ One by One, fill=orchestLight, text width= 6.5em, anchor=center, align=center, font=[ WebLINX Lù et al. (2024), Co-STORM Jiang et al. (2024), Agency Task Sharma et al. (2024),
Help Feedback Mehta et al. (2024), Ask Before Plan Zhang et al. (2024c), leaf, fill=orchestLight, text width=42.2em, anchor=center, align=center, text centered, font = ] ] [ Simultaneous, fill=orchestLight, text width= 6.5em, anchor=center, align=center, font=[ MTOM Zhang et al. (2024b), HRT-ML Liu et al. (2024b), DPT Agent Zhang et al. (2025),
CoELA Zhang et al. (2024a), leaf, fill=orchestLight, text width=42.2em, anchor=center, align=center, text centered, font = ] ] ] [ Temporal
Synchronization, fill=orchestLight, text width= 7.2em, anchor=center, align=center, font=[ Synchronous, fill=orchestLight, text width= 6.5em, anchor=center, align=center, font=[ AgentCoord Pan et al. (2024a), HRC Assembly Gkournelos et al. (2024),
HRC Manufa Lim et al. (2024), leaf, fill=orchestLight, text width=42.2em, anchor=center, align=center, text centered, font = ] ] [ Asynchronous, fill=orchestLight, text width= 6.5em, anchor=center, align=center, font=[ Collaborative Gym  Shao et al. (2024), MetaGPT Hong et al. (2023), TaPA Wu et al. (2023), leaf, fill=orchestLight, text width=42.2em, anchor=center, align=center, text centered, font = ] ] ] ] [ Communication
(Sec 3.5), fill=communClr, text width= 7em, anchor=center, align=center, font=[ Structure, fill=communLight, text width= 5em, anchor=center, align=center, font=[ Centralized, fill=communLight, text width= 6em, anchor=center, align=center, font=[ Drive As You Speack Cui et al. (2024), Organized Teams Guo et al. (2024b),
HRC Manufa Lim et al. (2024), DigiRL Bai et al. (2024), leaf, fill=communLight, text width=45em, anchor=center, align=center, text centered, font = ] ] [ Decentralized, fill=communLight, text width= 6em, anchor=center, align=center, font=[ Prison Dilemm Jiang et al. (2025), InteractGen Sun et al. (2024b), TaPA Wu et al. (2023),
MINT Wang et al. (2024b), leaf, fill=communLight, text width=45em, anchor=center, align=center, text centered, font = ] ] [ Hierarchical, fill=communLight, text width= 6em, anchor=center, align=center, font=[ FineArena Xu et al. (2025), AgentCoord  Pan et al. (2024a),
EasyLAN Pan et al. (2024b), Hierarchical Agent Liu et al. (2023b), EmoAgent Qiu et al. (2025), leaf, fill=communLight, text width=45em, anchor=center, align=center, text centered, font = ] ] ] [ Mode, fill=communLight, text width= 5em, anchor=center, align=center, font=[ Conversation, fill=communLight, text width= 6em, anchor=center, align=center, font=[ DigiRL Bai et al. (2024), WebLINX Lù et al. (2024), Organized Teams Guo et al. (2024b),
ConvCodeWorld Han et al. (2025), Help Feedback Mehta et al. (2024), leaf, fill=communLight, text width=45em, anchor=center, align=center, text centered, font = ] ] [ Observation, fill=communLight, text width= 6em, anchor=center, align=center, font=[ AXIS Lu et al. (2024), Attentive Supp Tanneberg et al. (2024a), EasyLAN Pan et al. (2024b),
MineWorld Guo et al. (2025), leaf, fill=communLight, text width=45em, anchor=center, align=center, text centered, font = ] ] [ Message
Pool, fill=communLight, text width= 6em, anchor=center, align=center, font=[ SMALL Wang et al. (2024c), HRT-ML Liu et al. (2024b), MetaGPT Hong et al. (2023), leaf, fill=communLight, text width=45em, anchor=center, align=center, text centered, font = ] ] ] ] ]
Figure 2: Taxonomy of LLM-based Human-Agent Systems. A more detailed and structured categorization of representative works is provided in the appendix (Table D.1 and D.2).
2 LLM-Based Human-Agent Systems
We define LLM-based human-agent systems as interactive frameworks where humans actively provide additional information, feedback, or control during interaction with an LLM-powered agent to enhance system performance, reliability, and safety Feng et al. (2024); Shao et al. (2024); Mehta et al. (2024). The core idea is synergy: combining unique human strengths—like intuition, creativity, expertise, ethical judgment, and adaptability—with LLM agent capabilities such as vast knowledge recall, computational speed, and sophisticated language processing. LLM-HAS builds upon core LLM agent components but places critical emphasis on the human’s interactive role and capabilities:
  1. (1)
    Providing Information / Clarification: Humans provide additional information that agents lack or cannot reliably infer, such as login credentials, payment details, domain expertise, constraints, or resolve ambiguities, helping agents interpret situations more accurately Naik et al. (2025); Kim et al. (2025).
  2. (2)
    Providing Feedback / Error Correction: Humans evaluate agent outputs and provide feedback, ranging from simple ratings to complex critiques, demonstrations or corrections, effectively guiding agents’ adjustment Gao et al. (2024b); Dutta et al. (2024); Li et al. (2024b).
  3. (3)
    Taking Control / Action: In high-stakes or sensitive scenarios (e.g., healthcare, privacy, or ethics), humans retain the authority to override, redirect, or halt agent actions, ensuring accountability, safety, and alignment with human values Chen et al. (2025); Natarajan et al. (2025); Xiao and Wang (2023).
Figure 1 provides a generalized overview of LLM-based human-agent systems. These systems operate within a defined Environment (e.g., physical world, simulation) that provides context and stimuli. Human & Agent Profiling characterizes the participants’ roles and goals, and the agent’s core LLM engine augmented with capabilities like planning, memory, and tool use. Human Feedback can occur during different phases in various types and granularities. Human-Agent Interaction Types may be collaborative (most common), competitive, cooperative, or mixed. The Orchestration layer governs high-level coordination, choosing a task strategy (e.g., sequential one-by-one versus parallel simultaneous execution) and a temporal synchronization mode (real-time synchronous exchanges versus delayed asynchronous workflows) so that each actor acts at the right moment. The Communication layer specifies how information flows, defining message structure (centralized, decentralized, hierarchical) and mode (conversation, observation signals, or shared message pools). The effective interplay and configuration of these components, along with various human feedback, are critical for tailoring the system to specific tasks and optimizing the overall system’s performance.
3 Core Components
In this section, we examine LLM-HAS through five core aspects: environment & profiling, human feedback, interaction type, orchestration paradigm, and communication. These dimensions provide a unified standard for analyzing existing work and guiding the design of future systems.
3.1 Environment and Profiling
Environment Setting. The environment in LLM-HAS defines a shared interaction space that can exist either in the physical world, such as offices Sun et al. (2024b), or in fully simulated virtual environments where agents and humans engage under controlled conditions Sun et al. (2024b); Zhang et al. (2024a); Guo et al. (2024b). These systems can be configured in various ways, including single-human single-agent, single-human multi-agent, multi-human single-agent, and multi-human multi-agent setups, each reflecting different collaboration dynamics and complexities.
Human & Agent Profiling. Human participants can be broadly categorized as lazy or informative users. Lazy users provide minimal guidance, typically offering evaluative feedback such as binary correctness or scalar rating. In contrast, informative users engage deeply by offering demonstrations, detailed guidance, refinements, or even taking over parts of the task Wang et al. (2024b); Liu et al. (2024b); Han et al. (2025). On the other side, agents are profiled by their roles and capabilities, which range from versatile general assistants to specialized experts in mathematics, engineering, medicine, or robotic cleaning, each adapted to the particular demands of its operational context Guo et al. (2024a); Samuel et al. (2024).
3.2 Human Feedback
Dimension Category Definition Summary Key Characteristics / Trade-offs Example Work
Type Evaluative User provides an assessment of the agent’s output quality, typically as binary assessment, scalar rating, or preference ranking. ① Easy to collect, scalable. ② Less specific signal for improvement. EmoAgent (Qiu et al., 2025), MINT (Wang et al., 2024b), SOTOPIA (Zhou et al., 2024)
Corrective User offers edits or fixes to the agent’s behavior. ① Highly informative, clear signal for improvement. ② Higher user effort, often fine-grained & interactive. SymbioticRAG (Sun et al., 2025), SWEET-RL (Zhou et al., 2025), AI Chains (Wu et al., 2022)
Guidance User proactively provides instructions, demonstrations, or critiques to shape the agent’s behavior. ① Bootstraps learning, conveys complex goals, proactive alignment. ② Requires clear specification from user. Drive As You Speack (Cui et al., 2024), Hierarchical Agent(Liu et al., 2023b), Ask Before Plan (Zhang et al., 2024c)
Implicit Inferred by the agent observing user actions or control signals, rather than explicitly stated or direct output modifications. ① Natural, unobtrusive collection. ② Ambiguous, requires careful interpretation. MTOM(Zhang et al., 2024b), Attentive Supp. (Tanneberg et al., 2024a), MineWorld (Guo et al., 2025)
Granularity Coarse-grained / Holistic Single assessment/signal for an entire agent output, trajectory, or task outcome. ①  Simple for user, good for overall assessment ②  Obscures specific errors, less precise learning signal. AssistantX (Sun et al., 2024a), Help Feedback (Mehta et al., 2024), AXIS (Lu et al., 2024)
Fine-grained / Segment-Level Feedback targeting specific parts of agent output, actions, or process. ①  Precise learning signal, crucial for debugging complex skills ②  Potentially higher user effort/burden. Collaborative Gym (Shao et al., 2024), Prison Dilemm (Jiang et al., 2025), FineArena (Xu et al., 2025)
Phase Initial Setup & Goal Definition Feedback provided before task execution, configuring the agent system and defining the task, goals, constraints, and preference. ①  Initial and proactive alignment, prevents costly errors, sets constraints ②  Requires upfront user input. AgentCoord (Pan et al., 2024a), GDfC (Wang et al., 2025c), SMALL (Wang et al., 2024c)
During Task Execution Online, interactive feedback while the agent is actively performing the task, enabling real-time adaptation. ①  Enables real-time adaptation, crucial for dynamic/collaborative tasks ②  Requires timely notification and responsive interfaces. InteractGen (Sun et al., 2024b), CowPilot (Huq et al., 2025), EasyLAN (Pan et al., 2024b)
Post-Task Eval. & Refinement Feedback provided after task completion to assess outcomes and provide suggestions for future use. ①  Non-disruptive, good for aggregate data/offline learning ②  No impact on completed task. HRT-ML (Liu et al., 2024b), M3HF (Wang et al., 2025d), MAIH (Wang et al., 2024c)
Table 1: Dimensions of Human Feedback in LLM-based human–agent systems, including feedback type, granularity, and phase. For each dimension, a summary, key characteristics, and example works are provided for comparison. A detailed overview of human feedback types and their subtypes is provided in our appendix (Table B.1).
Human Feedback Type. We categorize human feedback as evaluative, corrective, guidance, and implicit feedback. (1) Evaluative Feedback provides an assessment of the agent’s output quality, typically as preference ranking, scalar rating, or binary assessment. A prime example is preference ranking, where users compare agent outputs, forming the basis of Reinforcement Learning from Human Feedback (RLHF) Chaudhari et al. (2024). Alternatively, platforms like Uni-RLHF Yuan et al. (2024) support scalar ratings or binary assessments. (2) Corrective Feedback offers direct edits or fixes to the agent’s behavior. For instance, the PRELUDE Gao et al. (2024a) framework learns latent preferences from user edits made to agent-generated text. (3) Guidance Feedback means the human proactively provides instructions, critiques, or demonstrations to shape the agent’s behavior. Agents like InteractGen Sun et al. (2024b), AutoManual Chen et al. (2024a) can be bootstrapped using initial demonstrations, while methods like Self-Refine Choudhury and Sodhi (2025) employ iterative critiques and refinements to improve outputs. (4) Implicit Feedback is inferred by the agent observing user actions or control signals, rather than explicitly stated or direct output modifications. For example, an agent might learn user priorities by observing how a user adjusts control sliders in a system like VeriPlan Lee et al. (2025a), or infer preferences by analyzing user behaviors like clicks and purchases in frameworks such as AgentA/B Wang et al. (2025a). This contrasts with corrective feedback, where the user directly edits the output; here, the agent interprets the user’s independent actions or control choices.
Human Feedback Granularity. Human feedback also varies in granularity, from coarse-grained, holistic judgments to fine-grained, segment-level critiques. (1) Coarse-grained/Holistic feedback provides a single assessment for the entire agent output. Standard RLHF often relies on holistic preferences between complete responses, which simplifies feedback collection but struggles with credit assignment in complex tasks. (2) Fine-grained/Segment-Level Feedback by contrast, targets specific parts (e.g., sentences, paragraphs, code blocks). This is crucial in environments like ConvCodeWorld Han et al. (2025), where feedback pertains to specific conversational turns or generated code segments, or in annotation tasks like PDFChatAnnotator Tang et al. (2024), where feedback applies to specific annotations or parts of the document. This finer granularity provides more precise learning signals, crucial for debugging complex behaviors.
Human Feedback Phase. Human feedback can be incorporated at different phases of the LLM-agent pipeline Wang et al. (2025d). (1) Initial Setup & Goal Definition occurs before task execution, configuring the agent system and defining goals, such as setting coordination strategies (AgentCoord Pan et al. (2024a)) or critiquing plans before execution (Ask-before-Plan Zhang et al. (2024c)). (2) During Task Execution involves online, interactive feedback while the agent is actively performing the task, enabling real-time adaptation. Examples include interactive instruction editing (InstructEdit Wang et al. (2023)), mid-task refinements (Mutual Theory of Mind Zhang et al. (2024b), Collaborative Gym Shao et al. (2024)), or online interventions (HG-DAgger Kelly et al. (2019)). (3) Post-Task Evaluation & Refinement happens after task completion to assess outcomes and provide feedback for future use. Frameworks like MAIH (Wang et al., 2024c) and EmoAgent (Qiu et al., 2025) apply feedback loops after initial generation for benchmarking or offline learning, while AdaPlanner Sun et al. (2023) archives successful plans post-task as skills for future use. Table 1 summarizes different dimensions of human feedback, key characteristics, and example work.
3.3 Human-Agent Interaction Types
Interaction types define how individuals communicate, exchange information, and take actions with one another. In LLM-HAS, interactions tend to be more dynamic and complex compared to multi-agent systems. This complexity arises from the various roles and responsibilities assigned to both human agents and those based on LLMs, necessitating a finer-grained framework to describe their collaborative behaviors. The following categorization highlights the three key interaction types: Collaboration, Competition, and Coopetition.
3.3.1 Collaboration
Collaborations are by far the most common interaction and foundational interaction, which involve humans and LLM-based agents working together to achieve a common goal. This partnership combines human creativity and contextual understanding with LLM-based agents to address challenges and improve the efficiency and quality of results (Vats et al., 2024; Du et al., 2024; Sun et al., 2025). Depending on the type of collaboration considered, it can be categorized into four main fine-grained subtypes (Figure 3): (1) Delegation & Direct Command (Kiewiet and McCubbins, 1991), (2) Supervision (Loganbill et al., 1982), (3) Cooperation (Rand and Nowak, 2013), and (4) Coordination (Turvey, 1990).
Figure 3: The subtype of the collaboration between humans and LLM-based agents.
Delegation & Direct Command. In this interaction modality, a controlling party, usually a human, assigns explicit tasks to the LLM-based agent by providing clear and direct instructions. The agent is expected to execute these directives autonomously or on behalf of humans, ensuring that responsibilities are well-defined and actions align with the system’s overarching objectives. Unlike supervision, where strategies can be dynamically adjusted in response to new situations, delegation involves providing instructions upfront. This means the agent follows a predetermined set of tasks rather than adapting to the situation. For instance, an investor specifies their risk preference to the agent executing the investment strategy like FineArena (Xu et al., 2025), or a driver utters the command to LLM-based agent like Drive as you Speak (Cui et al., 2024).
Supervision. Supervision is the process by which one party, usually a human operator, oversees, monitors, and guides the actions of an LLM-based agent. This involves real time evaluation and intervention to ensure the agent’s output aligns with established goals and quality standards. Supervision also encompasses setting alert thresholds and providing corrective inputs when deviations occur. By maintaining a continuous feedback loop between the human and the agent, supervision helps calibrate agent behavior, catch and mitigate errors before they propagate and build confidence in the system. It also enables agents to handle routine tasks with increasing independence. For instance, agents notify humans to verify alignment (Liu et al., 2023b), and teleoperators monitor the LLM-generated motion plans (Liu et al., 2023a).
Cooperation. Cooperation refers to the voluntary and joint efforts of multiple parties to achieve agreed-upon goals. This collaboration type combines the various efforts and outcomes of different individuals and LLM-based agents toward a common objective. It emphasizes collective commitment, mutual assistance, and the pooling of resources to attain a shared result, thereby fostering a collaborative problem-solving environment. For instance, the human robot coordination in household activities Chang et al. (2024), the cooperative embodied language agent (CoELA) (Zhang et al., 2024a), human designers collaborating with an LLM-based agent (Sharma et al., 2024).
Coordination. Coordination is the organized process of aligning and synchronizing the actions of humans and LLM-based agents to achieve a shared objective. Unlike cooperation, the key idea behind coordination is to avoid conflict and bias in both humans and LLM-based agents to reach the final goal. It involves clear communication, strategic planning, and the intentional division of tasks, ensuring that individual efforts are harmonized and effectively integrated to support common goals. For example, humans and agents work in a shared workspace to complete interdependent tasks (Zhang et al., 2024b), human–agent integration supports adaptive decision-making (Sun et al., 2024b), and the collaborative framework facilitates coordination between humans and agents (Pan et al., 2024a).
3.3.2 Competition
Competition is a form of interaction where participants aim to achieve their own goals, which often conflict with the objectives of others. In the LLM-HAS, competition emerges when agents or humans seek to enhance their personal performance or obtain resources, even if it negatively impacts collective results. In addition, competition also necessitates effective balancing mechanisms, like performance regulation or conflict resolution strategies, to prevent unproductive behaviors and ensure that the overall goals of the system remain intact. For instance, the SOTOPIA framework simulates social behaviors between humans and LLM-based agents (Zhou et al., 2024).
3.3.3 Coopetition
Coopetition is an interaction where cooperation and competition coexist at the same time. Within this interaction, participants collaborate on shared tasks or mutual goals while also seeking to outdo each other to improve their own performance or gain extra advantages. In terms of the LLM-HAS, this dual aspect implies that agents and human may join forces to address complex issues while competing in specific domains such as efficiency or precision. This approach not only combines the strengths of both collaboration and competition, but also fosters innovation driven by competitive incentives while also reaping the benefits of cooperative synergy. Successfully managing coopetition typically requires mechanisms for building trust and adaptable strategies that reconcile collective advantages with personal aspirations, which is a challenge for the LLM-HAS. For example, humans and agents play the prisoner’s dilemma in the shared workspace (Jiang et al., 2025).
3.4 Orchestration Paradigm
The orchestration paradigm in LLM-HAS refers to how tasks and interactions are managed between humans and agents, covering two dimensions in our survey: Task Strategy (ordering) and Temporal Synchronization (timing). Table 2 summarizes the two dimensions of the orchestration paradigm.
3.4.1 Task Strategy
In LLM-HAS, the chosen task strategy, defined by the order and grouping of tasks performed by humans and agents, significantly impacts task execution effectiveness and efficiency. These strategies can typically be categorized into one-by-one and simultaneous paradigms.
Orchestration Paradigm Description
Task Strategy What order and grouping of tasks do participants perform?
One-by-One Actors take turns (e.g., human plans $\rightarrow$ agent executes $\rightarrow$ human reviews $\rightarrow$ agent refines).
Simultaneous Actors work in parallel (e.g., agent streams partial suggestions while human types).
Temporal Synchronization When and how tightly do actors’ steps need to align in time?
Synchronous (1) Real-time interaction: Humans and agents communicate simultaneously or in immediate sequence; (2) Immediate response: Participants expect or require prompt feedback. (e.g., live chat session, real-time voice assistant).
Asynchronous (1) Delayed interaction: Communication occurs without the expectation of immediate responses; (2) Flexible timing: Participants can respond at their convenience. (e.g., email queues, human leaves comments, agent processes offline).
Table 2: Orchestration paradigms in LLM-based human–agent systems encompass two orthogonal dimensions: task strategy, which can be one-by-one or simultaneous, and temporal synchronization, which can be synchronous or asynchronous.
One-by-One. The one-by-one strategy requires participants (humans and LLM-based agents) to perform tasks sequentially, taking clearly defined turns. For example, a human first outlines a plan, the agent then executes the task, the human subsequently reviews the output, and finally, the agent refines its work based on feedback Liu et al. (2024a); Zhou et al. (2025). Such sequential interaction helps maintain a clear order of execution and reduces the complexity associated with concurrent task management. However, this rigidity may limit overall efficiency and flexibility, especially in dynamic scenarios requiring parallel processing or rapid interaction cycles Bansal et al. (2024); Guo et al. (2024b).
Simultaneous. Simultaneous strategy describes an interaction pattern in which LLM-based agents and humans respond concurrently in real time. Compared to the one-by-one strategy, the simultaneous approach more closely mirrors real-world conditions encountered in many simulation tasks Wang et al. (2025d); Zhang et al. (2025). However, this strategy demands sophisticated mechanisms to handle latency mitigation and seamless coordination between participants.
3.4.2 Temporal Synchronization
Temporal synchronization in LLM-HAS refers to the timing and coordination of interactions between humans and agents. It significantly influences system responsiveness, user experience, and task efficiency. It can be broadly categorized into two modes: synchronous and asynchronous.
Synchronous. Synchronous interaction involves real-time interactions where humans and agents engage simultaneously. Immediate response is expected, facilitating dynamic exchanges. Examples include live chat sessions, real-time voice assistants (e.g., Siri, Alexa), and collaborative decision-making scenarios Zhang et al. (2024b); Liu et al. (2023b). This mode is advantageous for tasks requiring rapid responses, immediate clarification, or real-time collaboration Mehta et al. (2024); Han et al. (2025).
Asynchronous. In contrast, asynchronous interaction occurs without the necessity for immediate responses. Participants can engage at their convenience, allowing for flexibility in communication. Examples include email exchanges, message queues, ticket-based support systems, and task assignments where agents process and report outcomes independently Shao et al. (2024); Zhang et al. (2025). Asynchronous communication is beneficial for complex issues that require thoughtful analysis or when participants are in different time zones Sun et al. (2024b, a).
3.5 Communication
In LLM-HAS, communication serves as the fundamental mechanism defining the transmission, reception, and transformation of information between humans and LLM-based agents. It focuses specifically on how information flows across participants to support effective interaction and mutual understanding. Unlike LLM-based multi-agent systems Yan et al. (2025), human-agent systems introduce a unique dimension (i.e., flexible, and cognitively diverse human participation). This leads to a broader and more complex communication landscape, encompassing both human-to-agent and agent-to-agent exchanges, each influenced by human interpretability, feedback style, and interaction latency.
To systematically analyze communication behavior in such systems, we propose a two-dimensional taxonomy that captures the communication behavior characteristics of humans and agents from macro-structures to micro-interaction rules. Specifically, we divide this section into the following parts: Communication Structure, which describes the macro-level organization of information channels, and Communication Mode, which characterizes the micro-level methods of message exchange.
3.5.1 Communication Structure
Communication structure refers to the organizational structure of agents, including both humans and agents, in LLM-HAS. It determines how information flows at the macro level and shapes the rules of interaction at the micro level. While originally developed for LLM-based multi-agent environments Guo et al. (2024a), these structures have been effectively adapted to human-agent scenarios by treating humans as specialized agents. In such systems, the communication structure not only governs the efficiency of information exchange but also significantly impacts the system’s adaptability, scalability, and robustness to human variability. We categorize the representative structures into three types: Centralized, Decentralized, and Hierarchical.
In Centralized structure, one primary agent or a group of core agents acts as a central node to coordinate all communications within the system. This central agent manages interactions among other agents, simplifying coordination and minimizing conflicts Cui et al. (2024). Decentralized structure employs peer-to-peer communication, enabling direct interactions among agents without centralized control. Agents autonomously manage their communications based on systemic information, enhancing system flexibility, adaptability, and robustness Shao et al. (2024); Driess et al. (2023). In addition, Hierarchical structure organizes agents into clearly defined levels, assigning distinct roles and responsibilities according to their position within the hierarchy Liu et al. (2023b); Pan et al. (2024b). High-level agents typically fulfill managerial or strategic roles, providing overarching guidance and supervision, while lower-level agents perform specialized tasks and execute detailed operations.
3.5.2 Communication Mode
Communication mode defines the manner through which humans and agents exchange information within LLM-HAS. Specifically, communication mode describes the methods employed by participants to transmit, acquire, and utilize information, critically shaping interaction efficiency and the overall performance of the system. Broadly, communication modes can be categorized into three primary approaches: Conversation, Observation, and Shared Message Pool.
Conversation. The conversation-based mode is currently the most prevalent and intuitive approach in LLM-HAS, wherein agents and humans directly engage through natural language dialogues. This interaction format typically utilizes conversational interfaces that allow iterative exchanges, questions, clarifications, and dynamic responses, facilitating efficient collaboration and mutual understanding Shao et al. (2024). For instance, conversational LLM agents can assist users by answering queries, explaining complex concepts, or collaboratively solving reasoning tasks through iterative dialogues Wang et al. (2024b). While intuitive and flexible, conversational interactions rely significantly on the communicative clarity and dialogue management capabilities of LLM agents.
Observation. In the observation-based communication mode, agents acquire information implicitly by observing participants behaviors, decisions, or interactions within their environment, rather than through explicit verbal communication. This mode leverages indirect signals, including user actions, feedback cues, or behavioral traces, to infer intentions, preferences, or states Lu et al. (2024). For example, an LLM-driven tutoring system may adaptively provide targeted instructions by continuously observing student problem-solving behaviors without explicit verbal queries Pan et al. (2024b). However, relying solely on observational signals can introduce ambiguity, potentially impacting inference accuracy unless complemented by robust inferential mechanisms.
Message Pool. The shared message pool mode involves agents and humans exchanging information through a common information repository. Participants publish messages or data into a message pool, subscribing and retrieving relevant messages based on specific interests or tasks Sun et al. (2024a). This approach significantly simplifies direct agent-to-agent or human-to-agent interactions, reduces communication complexity, and enhances information management efficiency. A prominent example includes the MetaGPT framework Hong et al. (2023), where LLM-based agents collaboratively retrieve information dynamically from a shared message pool, streamlining cooperation and information dissemination. Despite these advantages, shared message pools must carefully manage access control to avoid information conflicts or inefficient retrieval.
Domain Datasets & Benchmarks Proposed or Used by Data Link
Embodied AI TaPA TaPA (Wu et al., 2023) Link
EmboInteract InteractGen (Sun et al., 2024b)
AssistantX AssistantX (Sun et al., 2024a)
IGLU Multi-Turn Help Feedback (Mehta et al., 2024) Link
PARTNR PARTNR (Chang et al., 2024) Link
MINT MINT (Wang et al., 2024b) Link
C-WAH REVECA (Seo et al., 2025) Link
HSRI HSRI (Lee et al., 2025b)
Conversational Systems WEBLINX WebLINX (Lù et al., 2024)
Ask-before-Plan Ask Before Plan (Zhang et al., 2024c) Link
Agency Dialogue Agency Task (Sharma et al., 2024)
WildSeek Co-STORM (Jiang et al., 2024) Link
MINT MINT (Wang et al., 2024b) Link
HOTPOTQA ReHAC (Feng et al., 2024) Link
StrategyQA ReHAC (Feng et al., 2024) Link
Software Development MINT MINT (Wang et al., 2024b) Link
InterCode ReHAC (Feng et al., 2024) Link
ColBench SWEET-RL (Zhou et al., 2025) Link
ConvCodeWorld ConvCodeWorld (Han et al., 2025) Link
ConvCodeBench ConvCodeWorld (Han et al., 2025) Link
Gaming CuisineWorld MindAgent (Gong et al., 2023) Link
MineWorld MineWorld (Guo et al., 2025) Link
Finance FinArena-Low-Cost FineArena (Xu et al., 2025) Link
Table 3: Datasets and Benchmarks across various domains.
4 Applications
A diverse range of applications has emerged for LLM-HAS. We elaborate on the five most frequent domains below and summarize corresponding datasets and benchmarks in Table 3. With new applications appearing almost weekly in this fast-growing field, we maintain an open-source GitHub repository to track recent developments.
Embodied AI. Applications in Embodied AI involve various aspects of dynamic and complex real-world tasks, benefiting from valuable human feedback and interactions in LLM-HAS. Ye et al. (2023) explores incorporating LLMs in human-robotic collaboration assembly tasks, allowing seamless communication between robots and humans and increasing trust in human operators. To address the challenges of false planning due to suboptimal environment changes, Seo et al. (2025) proposes REVECA to enable efficient memory management and optimal planning. Additionally, Tanneberg et al. (2024b) extends the agents’ collaboration with a group of humans via Attentive Support, enabling agents’ ability to remain silent to not disturb the group if desired.
Software Development. The inherently collaborative nature of software development makes human-agent collaboration vital to improve development efficiency Lu et al. (2024); Han et al. (2025); Zhou et al. (2025). Feng et al. (2024) introduces ReHAC framework, wherein agents are trained to determine the optimal stages for human intervention within the problem-solving process, offering improved generalizability over the traditional heuristic-based approaches. Building on this direction, Zhou et al. (2025); Han et al. (2025); Wang et al. (2024b) investigate a broader spectrum of human feedback types via multi-turn human-agent interactions. These approaches incorporate carefully designed optimization objectives to effectively capture more diverse and nuanced interactions between humans and agents.
Conversational Systems. In conversational systems, due to the frequent presence of ambiguity and lack of necessary information that agents cannot reliably infer, such as login credentials and payment details, effective human-agent collaboration constitutes a critical component of the system. Zhang et al. (2024c) introduces Proactive Agent Planning, wherein agents are trained to predict classification needs based on the user-agent conversational interactions and current environment, thereby leading to improved reasoning efficacy. Wu et al. (2022) introduces Chaining the LLM to improve the quality of task outcomes and enhance the transparency and controllability of the conversational systems.
Gaming. LLM-HAS are naturally well-suited to simulated gaming environments due to their dynamicity and sophistication. Proper human-agent interactions have been shown to enhance humans’ experience, satisfaction and understanding of both the environment and agents Gong et al. (2023); Gao et al. (2024c). Collaborative interactions also contribute to improved agents’ task performance and decision-making capabilities. For instance, MindAgent framework Gong et al. (2023) illustrates the efficacy of human-agent collaboration through measurable improvements in task outcomes when humans and agents work together. Mehta et al. (2024) demonstrates agents achieve improved outcomes when interacting with humans via autonomous confusion detection and clarification questions and inquiries. Ait et al. (2024) introduces Meta-Command Communication-based framework to enable effective human-agent collaboration. To address challenges related to execution latency while maintaining strong reasoning capabilities, Liu et al. (2023a) proposes Hierarchical Language Agent that promotes faster responses, stronger cooperation, and more consistent language communications.
Finance. Given the complexity of stock markets and financial systems, where investors’ strategies and risk preferences are critical determinants of outcomes, human-agent collaboration is increasingly recognized as a valuable paradigm. FinArena Xu et al. (2025) demonstrates the potential of integrating experienced investors with advanced AI agents to support stock prediction tasks. This collaborative framework has been shown to improve investment performance, yielding competitive annualized returns and Sharpe ratios Xu et al. (2025).
5 Implementation Tools and Resources
5.1 Human-Agent Framework
This section provides a detailed introduction to three representative open-source LLM-HAS frameworks: Collaborative Gym Shao et al. (2024), COWPILOT Huq et al. (2025), and DPT-Agent Zhang et al. (2025). They differ in key configuration aspects, including environment settings, interaction types, orchestration paradigms, and communication strategies. Specifically, Collaborative Gym Shao et al. (2024) facilitates asynchronous interactions among humans, agents, and task environments, supporting various simulated and real-world tasks such as travel planning, data analysis, and academic writing. It emphasizes flexible, real-time collaboration and evaluates both outcomes and interaction quality, making it a robust tool for studying human-agent dynamics. COWPILOT Huq et al. (2025) provides a framework for human-agent collaborative web navigation through a Chrome extension. It employs a "Suggest-then-Execute" model under human supervision, allowing dynamic interventions to enhance task completion rates and reduce human workload. It effectively demonstrates how human intervention can significantly improve agent performance. DPT-Agent Zhang et al. (2025) applies Dual Process Theory (DPT) to enable real-time simultaneous human-agent interactions. It features intuitive, fast decision-making and deliberative reasoning components, employing Theory of Mind and asynchronous reflection to manage latency and adapt dynamically to human actions. This approach excels in environments requiring immediate and adaptive responses.
Other frameworks, such as A2C Tariq et al. (2025), FinArena Xu et al. (2025), and human-robot collaboration framework Liu et al. (2023a), also contribute significantly to specific domains like cybersecurity, financial forecasting, and robotic manipulation, respectively. These frameworks further demonstrate the diverse potential and adaptability of LLM-HAS.
5.2 Datasets and Benchmarks
We summarize the commonly used datasets and benchmarks for Large Language Model-based Human-Agent Systems in Table 3. Diverse domains employ distinct methodologies for evaluating these systems, aligned closely with their unique application contexts. Within the domain of embodied AI, the primary approach involves simulated environments Sun et al. (2024b, a); Mehta et al. (2024), designed to assess how effectively agents cooperate and execute tasks in dynamic, interactive scenarios. Another significant domain, Conversational Systems, encompasses applications such as question answering Feng et al. (2024), website navigation Lù et al. (2024); Levy et al. (2024), design decision assistance Sharma et al. (2024), and travel planning Zhang et al. (2024c), adopting benchmarks that evaluate the ability of language models to function as user-aligned conversational assistants, ensuring interactions meet user expectations. Despite the extensive application coverage of current benchmarks, there remains a clear necessity for the development of more comprehensive and standardized benchmarking frameworks.
6 Challenges and Opportunities
In this section, we highlight some existing challenges and opportunities for LLM-HAS.
Human Flexibility and Variability. Human feedback varies widely in terms of role, timing, and style across various LLM-HAS. Humans are often subjective, influenced by their personalities, which means different individuals interacting with an LLM-HAS may lead to different outcomes and conclusions. This highlights the need and opportunity for i) thorough investigations or benchmarks on how varied human feedback affects entire systems, and ii) flexible frameworks that can support and adapt to diverse human feedback. In addition, humans, regarded as a “special agent” in the LLM-HAS, are subject to fewer restrictions and evaluations than LLM-based agents. This limits how the LLM-HAS can be improved because the impedance may be on the human side instead of the agent. This concern remains and requires a refined strategy to define the strict, fine interaction rule and evaluation equally for both human and LLM-based agents. Also, many studies today substitute real human participants with LLM simulated human proxies, failing to capture human input’s variety and unpredictability. The performance gap between humans and the simulated human remains unknown, potentially making the comparison incomparable.
Mostly Agent-Centered Work. In most LLM-HAS studies, guidance flows in a single direction, with humans evaluating agent outputs and providing corrective or evaluative feedback. Namely, the current studies are mostly agent-centered. However, enabling agents to observe human actions, detect errors or inefficiencies, and offer timely suggestions can transform collaboration and reduce human effort by leveraging agent intelligence. When agents act as instructors by proposing alternative strategies, drawing attention to overlooked risks, and reinforcing effective practices as tasks unfold in real time, both humans and agents benefit. We believe that exploring human-centered LLM-HAS, or shifting toward an equalized LLM-HAS, will unlock the full promise of teamwork between humans and agents.
Inadequate Evaluation Methodologies. In existing evaluation frameworks for LLM-HAS, improvements focus primarily on agent accuracy and static benchmarks, which ignore the real burden placed on human collaborators Ma et al. (2025). People dedicate varying amounts of time, attention and cognitive effort depending on the type and frequency of feedback they must provide, yet no standard metric captures this human workload or its impact on overall efficiency. Evaluation methods should measure factors such as time spent offering feedback, perceived mental workload and effort required to detect and correct errors, and they should cover every phase of the human agent collaboration from initial task assignment through post execution review. As human expertise and LLM-based agent capabilities merge to deliver unprecedented performance, both uncertainty and variability grow. A new evaluation approach or set of metrics that systematically and comprehensively quantifies contributions and costs for both humans and agents is essential to ensure truly efficient collaboration.
Unresolved Safety Vulnerabilities. Most LLM-HAS works emphasize improving agent performance and have left safety, robustness and privacy underexplored in the context of human interaction Qiu et al. (2025). As people and LLM-based agents collaborate in dynamic workflows, the risk of misaligned behavior, unexpected failures, or unintended disclosure of sensitive information grows. Humans engaging with these systems need clear safeguards around data sharing, error recovery protocols when agents behave unpredictably and privacy protections that cover every stage of the interaction. Robustness measures must ensure agents handle ambiguous or adversarial inputs without passing harm on to their human partners Glickman and Sharot (2025). Without studies that emphasize human experience in safety and privacy design, real-world deployments will struggle to gain trust or meet acceptable risk thresholds. Rigorous investigation of how safety, robustness and privacy shape human agent workflows from design through deployment is essential to build collaborations that are both effective and respectful of human needs.
Applications and Beyond. The potential of LLM-HAS extends well beyond current applications. Many opportunities remain to be explored in challenging domains such as healthcare, finance, scientific research, education, and so on. While fully autonomous LLM-based agent systems encounter difficulties in handling complex, long-term tasks and earning full trust in safety and reliability, the involvement of humans to provide additional information, feedback, and control allows LLM-HAS to greatly improve overall system performance and safety. This opens the door to impactful applications across a broad range of critical fields.
7 Conclusion
This paper presents a comprehensive review of LLM-based Human-Agent Systems. We introduce a structured taxonomy covering five core dimensions: environment and profiling, human feedback, interaction types, orchestration paradigms, and communication, and use it to classify and analyze existing research on LLM-HAS. We also summarize representative implementation frameworks, benchmark datasets, and evaluation metrics to support reproducibility and comparative analysis. Finally, we identify key challenges and unresolved issues in current LLM-HAS research. These issues remain major obstacles to the development of effective, adaptive, safe and trustworthy human-agent systems. We hope this review offers a comprehensive understanding of the LLM-HAS landscape and serves as a practical guide for future research.
Limitations
Although we strive to include a wide range of representative works (e.g., ACL, EMNLP, NAACL, EACL, COLM, NeurIPS, ICLR, ICML, etc.), some relevant research may not be included, especially recent preprints or interdisciplinary research in fields such as cognitive science. At the same time, although this review briefly discusses safety issues, it does not fully explore broader ethical and social impacts, including the allocation of responsibilities, long-term coexistence of humans and machines, and the consistency of values. These issues deserve further investigation in future work.
Appendix A Difference with Multi-Agent Systems
While both LLM-HAS and MAS involve collaboration among multiple entities, the key distinction lies in the nature and role of the collaborating parties Feng et al. (2024); Shao et al. (2024). Multi-agent systems are typically composed exclusively of autonomous agents—each designed to make decisions, communicate, and coordinate tasks with one another. In these MAS, each agent operates based on its own set of objectives and algorithms, and the overall behavior emerges from their interactions Tran et al. (2025); Guo et al. (2024a).
In contrast, LLM-based human–agent systems explicitly incorporate humans as active participants within the decision-making loop Feng et al. (2024). Rather than letting the system run purely on the combined strategies of several LLM-powered agents, these systems are engineered with mechanisms to allow human supervision, intervention, and feedback Mehta et al. (2024). This human-in-the-loop design is critical when balancing the strengths of LLMs—such as processing vast amounts of knowledge and performing rapid reasoning—with the need for contextual, ethical, and domain-specific judgments that humans uniquely provide Vats et al. (2024).
Furthermore, multi-agent systems often assume that the collaboration among agents can lead to a form of “collective intelligence” where agents work toward shared objectives Sun et al. (2024b). In many such frameworks, the communication protocols, coordination strategies, and role dynamics are all defined among non-human entities. In contrast, in human–agent systems, the interaction protocols are designed to enhance transparency and provide control for human decision-makers Shao et al. (2024). The system can selectively escalate issues for human review, enable corrective actions when the automated decision may be off-mark, and integrate human feedback to iteratively improve the agent’s performance over time Mehta et al. (2024).
Appendix B Human Feedback Type and Subtype
In this appendix, we present a detailed overview of human feedback types and their subtypes, as summarized in Table B.1. This table provides concise definitions and illustrates how humans provide feedback to LLM-based agents in LLM-HAS. While the main paper introduced the broad categories of evaluative, corrective, guidance, and implicit feedback, here we expand each category into more granular subtypes, ranging from scalar ratings and preference rankings to direct edits, demonstrations, and inferred behavioral signals. Recognizing these subtypes clarifies the ways in which humans interact with LLM agents, by offering precise instructions and well-defined tasks, to enhance the accuracy and quality of generated outputs. This deeper understanding empowers users to optimize their interactions with LLM-based agents. Additionally, the systematic breakdown of human feedback provides a foundation for cross-study comparisons. It underscores the diverse strategies through which human users can guide, correct, or collaborate with LLM-based agents in a more detailed way.
Human Feedback Type Description How it Helps Agents
Evaluative Feedback User provides an assessment of the agent’s output quality. Signals overall correctness or preference, guiding general alignment.
Preference Ranking User compares two or more agent outputs and selects the preferred one. Helps the agent learn relative quality and subjective nuances.
Scalar Rating User assigns a numerical score (e.g., 1–5) to the agent’s output. Provides a quantitative measure of satisfaction or quality.
Binary Assessment User indicates simple correctness (e.g., yes/no, thumbs up/down). Offers a basic signal of success or failure.
Corrective Feedback User modifies or directly improves the agent’s output. Provides explicit examples of desired output, enabling direct learning from errors.
Direct Edits / Refinements User manually changes the agent’s generated text or code. Shows the agent the precise correction needed.
Guidance Feedback User provides instructions or explanations to steer the agent. Offers deeper context, reasoning, or demonstrations for learning complex behaviors.
Demonstrations User shows the agent how to perform a task correctly. Teaches specific procedures or desired interaction patterns.
Instructions / Critiques User provides natural language explanations, critiques, or step-by-step guidance. Helps the agent understand why an output is wrong and how to improve.
Implicit Feedback Agent infers user preference from their behavior. Reveals preferences and usability issues without explicit feedback requests.
Human Action / Control Human directly takes actions and control. Collaborates with humans to effectively finish tasks or learns from human actions.
Table B.1: Human Feedback Types and Subtypes. The subtypes of evaluative feedback includes preference ranking, scalar rating, and binary assessment. The subtypes of corrective feedback includes the direct edits or refinement. The subtypes of guidance feedback includes the demonstration and instructions or critiques. The subtypes of implicit feedback include the human action or control.
Appendix C Evaluation Metrics
In this section, we introduce evaluation metrics specifically designed for human-agent systems across four key aspects: feedback mechanisms, adaptability, trust and safety, and communication methods. To evaluate feedback mechanisms, Liu et al. (2024b) assesses a human-robot teaming framework using multi-modal language feedback at varying frequency levels (inactive, passive, active, superactive). Metz et al. (2024) proposes seven metrics, expressiveness, ease, definiteness, context independence, precision, unbiasedness, and informativeness, to evaluate feedback quality. In the education domain, Seßler et al. (2025) adopts six dimensions based on educational feedback theory. Spencer et al. (2020) evaluates the Expert Intervention Learning (EIL) method by comparing robot performance with and without expert intervention. For adaptability, Hauptman et al. (2023) examines how human-LLM agents respond to cyber incidents under different levels of autonomy across five NIST-defined phases. For trust and safety, Levy et al. (2024) introduces a benchmark that evaluates web agents on their ability to comply with policies, avoid unsafe behavior, respect security constraints, and handle errors gracefully, including seeking user input when needed. Finally, Karten et al. (2023) assess four categories of communication methods in human-agent teaming, focusing on effectiveness and interpretability within simulated environments of Predator-PreyLowe et al. (2017) and Traffic JunctionSingh et al. (2018).
In addition to these aspects, AXIS Lu et al. (2024) and SYNERGAI Chen et al. (2024b) evaluate the effectiveness and robustness of human–LLM agent systems in the domains of operating systems and embedded AI, respectively. These studies highlight how evaluation criteria can vary significantly depending on the specific task or application context, reflecting differences in system constraints, performance expectations, and interaction complexity.
Appendix D Tables
Table D.1 catalogs the environmental configuration and human feedback type, and Table D.2 categorizes the interaction, orchestration, and communication of the current works, respectively.
Table D.1: ① Environment Configuration and ② Human Feedback to LLM-based agents in human–agent systems. Environment Configuration specifies whether a single or multiple humans collaborate with one or more LLM-based agents, while Human Feedback characterizes the type, subtype, granularity, and interaction phase of the human feedback to the LLM-based agents.
Environment Configuration Human Feedback to LLM-based Agent
Paper Venue Code/ Data Human LLM Agent Type Subtype Granularity Phase
Collaborative Gym (Shao et al., 2024) Arxiv’24 Link Single Single Corrective, Guidance Refinements, Instructions Segment During Task
MTOM (Zhang et al., 2024b) Arxiv’24 Single Single Implicit Human Action Segment During Task
FineArena (Xu et al., 2025) Arxiv’25 Single Multiple Guidance Demonstrations Segment, Holistic Initial Setup, During Task
Prison Dilemm (Jiang et al., 2025) Arxiv’25 Single Single Implicit Human Action Segment During Task
InteractGen (Sun et al., 2024b) THU’24 Multiple Multiple Guidance Demonstrations Segment Initial Setup, During Task
AI Chains (Wu et al., 2022) CHI’24 Single Single Corrective Refinements Segment During Task
Drive As You Speak (Cui et al., 2024) WACV’24 Single Single Guidance Demonstrations Holistic Initial Setup
AgentCoord (Pan et al., 2024a) Arxiv’24 Link Single Multiple Evaluative, Corrective Preference Ranking, Refinements Segment, Holistic Initial Setup, During Task
CowPilot (Huq et al., 2025) Arxiv’25 Link Single Single Corrective, Evaluative Binary Assessment, Refinements Segment During Task
EasyLAN (Pan et al., 2024b) Arxiv’24 Single Multiple Corrective, Guidance Demonstrations, Refinements Segment, Holistic During Task
Hierarchical Agent (Liu et al., 2023b) AAMAS’24 Single Multiple Guidance Demonstrations Segment During Task
SWEET-RL (Zhou et al., 2025) Arxiv’25 Link Single Single Corrective, Implicit Refinements, Human Action Segment Initial Setup, During Task
HRC Assembly (Gkournelos et al., 2024) CIRP’24 Single Multiple Guidance Demonstrations Segment During Task
REVECA (Seo et al., 2025) Arxiv’24 Single Multiple Guidance Demonstrations Holistic Initial Setup
AssistantX (Sun et al., 2024a) Arxiv’24 Link Multiple Multiple Implicit, Guidance Human Action, Demonstrations Holistic, Segment Initial Setup, During Task
MINT (Wang et al., 2024b) ICLR’24 Link Multiple Single Evaluative, Corrective, Guidance Binary Assessment, Refinements, Instructions Holistic During Task
Help Feedback (Mehta et al., 2024) EACL’24 Single Single Evaluative, Guidance Demonstrations, Instructions, Binary Assessment Holistic, Segment During Task
ConvCodeWorld (Han et al., 2025) ICLR’25 Link Single Single Guidance, Evaluative Demonstrations, Instructions, Binary Assessment Segment, Holistic During Task
ReHAC (Feng et al., 2024) ACL’24 Link Single Single Corrective Refinements Segment During Task
DPT Agent (Zhang et al., 2025) Arxiv’25 Link Single Single Guidance Instructions Holistic During Task
HRC Manipulation (Liu et al., 2023a) IEEE’23 Single Single Corrective, Guidance Demonstrations, Refinements Segment During Task
HRC DMP (Liu et al., 2024a) IEEE’24 Single Single Corrective, Guidance Refinements, Demonstrations Segment During Task
PARTNR (Chang et al., 2024) ICLR’25 Link Single Single Guidance Demonstrations Holistic Inital Setup
Organized Teams (Guo et al., 2024b) Arxiv’24 Link Single Multiple Guidance Demonstrations Holistic, Segment Initial Setup, During Task
CoELA (Zhang et al., 2024a) ICLR’23 Single Multiple Guidance Demonstrations Holistic, Segment Initial Setup, During Task
Agency Task (Sharma et al., 2024) EACL’24 Link Single Single Guidance Demonstrations Segment During Task
GDfC (Wang et al., 2025c) SME’25 Single Multiple Guidance, Evaluative Demonstrations, Binary Assessment, Preference Ranking Holistic, Segment Initial Setup, During Task, Post Task
PDFChatAnnotator (Tang et al., 2024) IUI’24 Single Single Corrective, Guidance Demonstrations, Refinements Segment During Task
Attentive Supp. (Tanneberg et al., 2024a) IEEE’24 Link Multiple Single Implicit, Guidance Demonstrations, Human Action Segment During Task
HRC Trust (Ye et al., 2023) IEEE’23 Single Single Guidance Demonstrations, Instructions Segment During Task
BPMN (Ait et al., 2024) Arxiv’24 Link Multiple Multiple Guidance, Corrective Instructions, Refinements Segment During Task, Post Task
Co-STORM (Jiang et al., 2024) EMNLP’24 Link Single Multiple Guidance Demonstrations Segment During Task
HRC Manufa. (Lim et al., 2024) IEEE’24 Single Single Corrective, Guidance Demonstrations, Refinements, Instructions Segment Initial Setup, During Task
A2C (Tariq et al., 2025) Arxiv’24 Link Multiple Multiple Guidance, Evaluative Binary Assessment, Instructions Holistic, Segment During Task
MindAgent (Gong et al., 2023) NAACL’24 Link Single Multiple Guidance Demonstrations Segment During Task
Ask Before Plan (Zhang et al., 2024c) EMNLP’24 Link Single Multiple Guidance Demonstrations Segment Initial Setup, During Task
SOTOPIA (Zhou et al., 2024) ICLR’24 Multiple Multiple Evaluative, Implicit Scaler Rating, Human Action Holistic, Segment During Task, Post Task
PaLM-(Driess et al., 2023) ICML’23 Link Single Single Guidance, Implicit Demonstrations, Human Action Holistic, Segment Initial Setup, During Task
TaPA (Wu et al., 2023) Arxiv’23 Link Single Single Guidance Demonstrations Holistic, Segment Initial Setup
MetaGPT (Hong et al., 2023) ICLR’24 Link Single Multiple Guidance Demonstrations Holistic Initial Setup
DigiRL (Bai et al., 2024) NeurIPS’24 Link Single Single Evaluative, Guidance Binary Assessment, Demonstrations Holistic During Task, Post Task
WebLINX (Lù et al., 2024) Arxiv’24 Link Single Multiple Guidance Demonstrations Holistic, Segment Initial Setup, During Task
MineWorld (Guo et al., 2025) Arxiv’25 Link Multiple Single Implicit Human Action Segment During Task
M3HF (Wang et al., 2025d) ICML’25 Multiple Multiple Evaluative, Guidance Binary Assessment, Instructions Segment, Holistic During Task, Post Task
SMALL (Wang et al., 2024c) Arxiv’24 Single Multiple Guidance Instructions Segment Initial Setup
MAIH (Wang et al., 2024c) Arxiv’24 Single Single Implicit Human Action Segment During Task, Post Task
HRT-ML (Liu et al., 2024b) Arxiv’24 Single Multiple Corrective, Guidance, Implicit Refinements, Instructions, Human Action Holistic Initial Setup, During Task, Post Task
AXIS (Lu et al., 2024) Arxiv’25 Single Multiple Evaluative, Implicit, Corrective Human Action, Refinements, Binary Assessment Holistic Initial Setup, During Task, Post Task
EmoAgent (Qiu et al., 2025) Arxiv’25 Single Multiple Corrective, Implicit, Guidance Human Action, Instructions, Binary Assessment Segment, Holistic During Task, Post Task
SymbioticRAG (Sun et al., 2025) Arxiv’25 Single Single Corrective, Implicit, Evaluative Binary Assessment, Refinements, Demonstrations, Instructions, Human Action Segment Initial Setup, During Task, Post Task
Table D.2: ① Interaction ② Orchestration ③ Communication in LLM-based human–agent systems. Interaction types capture the human and agent collaboration type; Orchestration covers task strategy and temporal synchronization; Communication describes how messages are structured and delivered in the system.
Interaction Orchestration Communication
Paper Venue Code/ Data Types Variant Strategy Synchronization Structure Mode
Collaborative Gym (Shao et al., 2024) Arxiv’24 Link Collaboration Cooperation, Delegation One-by-One Asynchronous Decentralized Conversation
MTOM (Zhang et al., 2024b) Arxiv’24 Collaboration Coordination, Cooperation Simultaneous Synchronous Decentralized Conversation
FineArena (Xu et al., 2025) Arxiv’25 Collaboration Delegation, Cooperation One-by-One Synchronous Hierarchical Conversation
Prison Dilemm (Jiang et al., 2025) Arxiv’25 Coopetition One-by-One Asynchronous Decentralized Conversation
InteractGen (Sun et al., 2024b) THU’24 Collaboration Cooperation, Delegation, Coordination One-by-One Asynchronous Decentralized Message Pool
AI Chains (Wu et al., 2022) CHI’24 Collaboration Cooperation One-by-One Synchronous Decentralized Conversation
Drive As You Speak (Cui et al., 2024) WACV’24 Collaboration Delegation One-by-One Synchronous Centralized Conversation
AgentCoord (Pan et al., 2024a) Arxiv’24 Link Collaboration Coordination One-by-One Synchronous Hierarchical Conversation
CowPilot (Huq et al., 2025) Arxiv’25 Link Collaboration Supervision, Delegation, Coordination One-by-One Synchronous Decentralized Conversation
EasyLAN (Pan et al., 2024b) Arxiv’24 Collaboration Delegation, Supervision One-by-One Synchronous Hierarchical Observation
Hierarchical Agent (Liu et al., 2023b) AAMAS’24 Collaboration Supervision, Delegation, Cooperation One-by-One Synchronous Hierarchical Conversation
SWEET-RL (Zhou et al., 2025) Arxiv’25 Link Collaboration Delegation One-by-One Synchronous Centralized Conversation
HRC Assembly (Gkournelos et al., 2024) CIRP’24 Collaboration Delegation, Cooperation One-by-One Synchronous Decentralized Conversation
REVECA (Seo et al., 2025) Arxiv’24 Collaboration Cooperation One-by-One Synchronous Decentralized Conversation
AssistantX (Sun et al., 2024a) Arxiv’24 Link Collaboration Delegation, Cooperation One-by-One Asynchronous Decentralized Message Pool
MINT (Wang et al., 2024b) ICLR’24 Link Collaboration Delegation, Cooperation One-by-One Synchronous Decentralized Conversation
Help Feedback (Mehta et al., 2024) EACL’24 Collaboration Supervision, Delegation, Cooperation One-by-One Asynchronous Decentralized Conversation
ConvCodeWorld (Han et al., 2025) ICLR’25 Link Collaboration Supervision, Delegation One-by-One Asynchronous Decentralized Conversation
ReHAC (Feng et al., 2024) ACL’24 Link Collaboration Coordination, Supervision One-by-One Synchronous Decentralized Conversation
DPT Agent (Zhang et al., 2025) Arxiv’25 Link Collaboration Coordination Simultaneous Asynchronous Decentralized Observation
HRC Manipulation (Liu et al., 2023a) IEEE’23 Collaboration Supervision, Delegation One-by-One Synchronous Decentralized Conversation
HRC DMP (Liu et al., 2024a) IEEE’24 Collaboration Delegation, Supervision One-by-One Synchronous Decentralized Conversation
PARTNR (Chang et al., 2024) ICLR’25 Link Collaboration Coordination, Cooperation Simultaneous Synchronous Decentralized, Centralized Observation
Organized Teams (Guo et al., 2024b) Arxiv’24 Link Collaboration Cooperation, Coordination One-by-One Synchronous Decentralized, Centralized, Hierarchical Conversation
CoELA (Zhang et al., 2024a) ICLR’23 Collaboration Cooperation, Coordination Simultaneous Synchronous Decentralized Conversation
Agency Task (Sharma et al., 2024) EACL’24 Link Collaboration Cooperation, Delegation One-by-One Synchronous Decentralized Conversation
GDfC (Wang et al., 2025c) SME’25 Collaboration Delegation One-by-One Synchronous Decentralized Conversation
PDFChatAnnotator (Tang et al., 2024) IUI’24 Collaboration Delegation One-by-One Synchronous Decentralized Conversation
Attentive Supp. (Tanneberg et al., 2024a) IEEE’24 Link Collaboration Coordination One-by-One Synchronous Decentralized Observation
HRC Trust (Ye et al., 2023) IEEE’23 Collaboration Delegation One-by-One Synchronous Decentralized Conversation
BPMN (Ait et al., 2024) Arxiv’24 Link Collaboration Coordination Simultaneous Asynchronous Decentralized Message Pool
Co-STORM (Jiang et al., 2024) EMNLP’24 Link Collaboration Coordination One-by-One Synchronous Centralized Conversation
HRC Manufa. (Lim et al., 2024) IEEE’24 Collaboration Delegation, Cooperation One-by-One Synchronous Centralized Conversation
A2C (Tariq et al., 2025) Arxiv’24 Link Collaboration Cooperation One-by-One Asynchronous Hierarchical Conversation
MindAgent (Gong et al., 2023) NAACL’24 Link Collaboration Coordination Simultaneous Synchronous Centralized Conversation
Ask Before Plan (Zhang et al., 2024c) EMNLP’24 Link Collaboration Coordination, Delegation One-by-One Synchronous Hierarchical Conversation
SOTOPIA (Zhou et al., 2024) ICLR’24 Collaboration, Competition, Coopetition Coordination, Cooperation One-by-One Synchronous Decentralized Conversation
PaLM-(Driess et al., 2023) ICML’23 Link Collaboration Delegation One-by-One Synchronous Decentralized Conversation
TaPA (Wu et al., 2023) Arxiv’23 Link Collaboration Delegation One-by-One Asynchronous Decentralized Conversation
MetaGPT (Hong et al., 2023) ICLR’24 Link Collaboration Coordination One-by-One Asynchronous Decentralized Message Pool
DigiRL (Bai et al., 2024) NeurIPS’24 Link Collaboration Delegation One-by-One Synchronous Centralized Conversation
WebLINX (Lù et al., 2024) Arxiv’24 Link Collaboration Delegation One-by-One Synchronous Hierarchical Conversation
MineWorld (Guo et al., 2025) Arxiv’25 Link Collaboration Delegation One-by-One Synchronous Decentralized Observation
M3HF (Wang et al., 2025d) ICML’25 Collaboration Cooperation One-by-One, Simultaneous Synchronous Centralized Message Pool
SMALL (Wang et al., 2024c) Arxiv’24 Collaboration Delegation One-by-One Asynchronous Hierarchical Message Pool
MAIH (Wang et al., 2024c) Arxiv’24 Collaboration Delegation, Cooperation, Coordination One-by-One, Simultaneous Asynchronous Decentralized, Hierarchical Message Pool
HRT-ML (Liu et al., 2024b) Arxiv’24 Collaboration Coordination, Cooperation One-by-One, Simultaneous Asynchronous, Synchronous Hierarchical, Centralized Message Pool, Conversation
AXIS (Lu et al., 2024) Arxiv’25 Collaboration Delegation One-by-One Synchronous Centralized Conversation, Observation
EmoAgent (Qiu et al., 2025) Arxiv’25 Collaboration Supervision, Coordination, Cooperation One-by-One Synchronous Hierarchical, Centralized Conversation, Observation
SymbioticRAG (Sun et al., 2025) Arxiv’25 Collaboration Cooperation, Supervision, Delegation One-by-One Synchronous Centralized Conversation